# DragText: Rethinking Text Embedding in Point-based Image Editing

### v1

@@ 8 Conclusion
In this study, we introduced DragText, a method emphasizing the critical role of text prompt in point-based image editing with diffusion models. We found that static text embeddings hinder the editing process by causing drag halting and loss of content. By optimizing text embeddings in parallel with image embeddings, DragText ensures that text and image embeddings remain strongly coupled, ensuring better content preservation and drag accuracy. Our approach integrates seamlessly with existing diffusion-based drag models and includes a text embedding regularization technique to retain semantic integrity, hence significantly improving the performance and fidelity of image editing.
Appendix A More Details on Prompt Engineering
@@ A.1 Implementation Details
$\mathbf{\hat{c}}^{k}=w^{k}\cdot\mathbf{\hat{c}}^{k}_{\text{org}}+(1-w^{k})\cdot\mathbf{\hat{c}}^{k}_{\text{int}}$
where $\mathbf{\hat{c}}^{k}_{\text{org}}$ is the original text embedding and $\mathbf{\hat{c}}^{k}_{\text{int}}$ is the intention text embedding. At the beginning of the point-based editing, $w^{k}=1$ so $\mathbf{\hat{c}}^{k}=\mathbf{\hat{c}}^{k}_{\text{org}}$. Conversely, as the dragging progresses and handle points approach target points, the weight value $w^{k}$ increases, resulting in a higher proportion of $\mathbf{\hat{c}}^{k}_{\text{int}}$. In this way, as the image is progressively edited, the proportion of the intention text embedding is gradually increased.
A.2 More Qualitative Results for Prompt Engineering
In Fig. 11, we present additional results for prompt engineering. Corresponding results for DragText are also presented to validate the effectiveness of our approach in comparison to prompt engineering. Prompt engineering is found to have little impact on alleviating drag halting. In contrast, DragText dragged handle points closer to target points, compared to the original text prompt, the intention text prompt, and their interpolation.
Figure 11: More qualitative results for prompt engineering (i.e. the intention text, and interpolate intention text). Red points and blue points imply handle points and their target points respectively. Overall, DragText moved the content at the handle point of the User Drag image closest to the target point. Additionally, in (b), it can be observed that interpolating the intention text with the original text makes it challenging to maintain the object’s style. In (d), DragText preserved the most semantics while dragging appropriately.
Appendix B More Details on DragText
@@
B.1 Pseudo Code of DragText
@@
Output: Output image $\mathbf{\hat{x}_{0}}$
@@
4:     $\mathcal{F}(\mathbf{\hat{z}}_{t}^{k},\mathbf{\hat{c}}^{k})\leftarrow U_{\theta}(\mathbf{\hat{z}}_{t}^{k};\mathbf{\hat{c}}^{k})$
5:     Update $\mathbf{\hat{z}}_{t}^{k}$ using motion supervision
6:     Update $\mathbf{\hat{c}}^{k}$ using text optimization
7:     $\mathbf{\hat{z}}_{t}^{k+1},\mathbf{\hat{c}}^{k+1}\leftarrow\mathbf{\hat{z}}_{t}^{k},\mathbf{\hat{c}}^{k}$
@@
9:$\mathbf{\hat{z}}_{t},\mathbf{\hat{c}}\leftarrow\mathbf{\hat{z}}_{t}^{K},\mathbf{\hat{c}}^{K}$
@@
11:     $\mathbf{\hat{z}}_{t-1}\leftarrow$ apply denoising to $\mathbf{\hat{z}}_{t}$ conditioned on $\mathbf{\hat{c}}$
12:$\mathbf{\hat{x}}_{0}\leftarrow\mathbf{\hat{z}}_{0}$
@@
B.2 Modifications for Integrating with Other Methods
@@ B.2.1 FreeDrag
$\small\mathcal{L}_{\text{ms}}=$ $\sum_{i=1}^{n}\sum_{q_{1}}\left\|\mathcal{F}_{q_{1}+d_{i}}(\mathbf{\hat{z}}_{t}^{k},\mathbf{\hat{c}}^{k})-\mathcal{T}_{i}^{k}\right\|_{1}$
$+\lambda_{\text{image}}\left\|\left(\mathbf{\hat{z}}_{t-1}^{k}-\text{sg}(\mathbf{\hat{z}}_{t-1}^{0})\right)\odot(\mathbbm{1}-M_{\text{image}})\right\|_{1}.$
@@
In FreeDrag, if the template feature is poorly learned (category (c)), the point not only reverts to its previous position but also reuses the template feature map without updating its values. Inspired by this approach, our DragText computes the text loss only during the (a) and (b) processes to align with the image. In cases categorized as (c), the text embedding is not used for optimization. Therefore, we define $\alpha_{i}$ as 1 for cases (1) and (2), and 0 for case (3). The text loss is then defined as follows:
$\small\mathcal{L}_{\text{text}}=$ $\sum_{i=1}^{n}\sum_{q_{1}}\alpha_{i}\left\|\mathcal{F}_{q_{1}+d_{i}}(\mathbf{\hat{z}}_{t}^{k},\mathbf{\hat{c}}^{k})-\mathcal{T}_{i}^{k}\right\|_{1}$
$+\lambda_{\text{text}}\left\|\left(\mathbf{\hat{c}}^{k}-\text{sg}(\mathbf{c}^{0}))\right)\odot M_{\text{text}}\right\|_{1}.$
@@
In DragText, during image optimization, $\mathbf{\hat{c}}^{k}$ does not undergo gradient descent and during text optimization the latent vector $\mathbf{\hat{z}}_{t}^{k}$ does not undergo gradient descent.
B.2.2 DragNoise
@@
The bottleneck features $\mathbf{s}_{t}$ effectively capture richer noise semantics and efficiently capture most semantics at an early timestep $t$. Thus, DragNoise optimizes the bottleneck feature $\mathbf{s}_{t}$ of the U-Net instead of the latent vector $\mathbf{z}_{t}$, thereby shortening the back-propagation chain. Accordingly, DragText optimizes $\mathbf{s}_{t}$ instead of $\mathbf{z}_{t}$ during the image optimization processes. For each iteration $k$, $\mathbf{\hat{s}}_{t}^{k}$ undergoes a gradient descent step to minimize $\mathcal{L}_{\text{ms}}$:
$\small\mathbf{\hat{s}}_{t}^{k+1}=\mathbf{\hat{s}}_{t}^{k}-\eta_{\text{ms}}\frac{\partial\mathcal{L}_{\text{ms}}(\mathbf{\hat{s}}_{t}^{k},\mathbf{\hat{c}}^{k})}{\partial\mathbf{\hat{s}}_{t}^{k}}.$
@@
In DragText, neither the latent vector $\mathbf{z}_{t}^{k}$ nor the bottleneck feature $\mathbf{s}_{t}^{k}$ undergoes gradient descent during the text optimization. Thus, the text optimization procedure is not modified in DragNoise.
B.2.3 GoodDrag
@@
GoodDrag alternates between dragging and denoising. This approach is different from traditional diffusion-based dragging methods. They generally execute all drag operations at once before denoising the optimized noisy latent vector $\mathbf{\hat{z}}_{t}$. During the denoising process, which involves sampling images $\mathbf{\hat{x}}_{0}$ from a noisy latent vector $\mathbf{\hat{z}}_{t}$, perturbations from dragging are corrected. However, if the denoising process is performed only after all drag operations are completed, the errors accumulate too significantly to be corrected with high fidelity. To address this, GoodDrag applies one denoising operation after $B$ image optimization and point tracking steps.
@@
For example, the latent vector $\mathbf{\hat{z}}^{k}_{t}$ has been denoised $\lfloor\frac{k}{B}\rfloor$ times, the drag optimization is performed at the timestep $t=T-\frac{k}{B}$. To ensure this process is consistent, the total number of drag steps $K$ should be divisible by $B$. Since DragText performs one text optimization step after one image optimization step, we sequentially repeat the image optimization, text optimization, and point tracking steps $B$ times, and then apply one denoising operation.
@@
$\small\mathcal{L}_{\text{ms}}=$ $\sum_{i=1}^{n}\sum_{q_{1}}\left\|\mathcal{F}_{q_{1}+4\cdot d_{i}}(\mathbf{\hat{z}}_{t}^{k},\mathbf{\hat{c}}^{k})-\text{sg}(\mathcal{F}_{q_{1}^{0}}(\mathbf{\hat{z}}_{t}^{0},\mathbf{\hat{c}}^{0}))\right\|_{1}$
$+\lambda_{\text{image}}\left\|\left(\mathbf{\hat{z}}_{t-1}^{k}-\text{sg}(\mathbf{\hat{z}}_{t-1}^{0})\right)\odot(\mathbbm{1}-M_{\text{image}})\right\|_{1},$
where $q_{1}=\Omega(h_{i}^{k},r_{1})$, and $q_{1}^{0}$ describes the square region centered at the original handle point $h_{i}^{0}$. Additionally, drag operations per denoising step $B=10$. Similarly, DragText ensures the handle point $h^{k}_{i}$ remains consistent with the original point $h^{0}_{i}$ during text optimization:
$\small\mathcal{L}_{\text{text}}=$ $\sum_{i=1}^{n}\sum_{q_{1}}\left\|\mathcal{F}_{q_{1}+4\cdot d_{i}}(\mathbf{\hat{z}}_{t}^{k},\mathbf{\hat{c}}^{k})-\text{sg}(\mathcal{F}_{q_{1}^{0}}(\mathbf{\hat{z}}_{t}^{0},\mathbf{\hat{c}}^{0}))\right\|_{1}$
$+\lambda_{\text{text}}\left\|\left(\mathbf{\hat{c}}^{k}-\text{sg}(\mathbf{c}^{0}))\right)\odot M_{\text{text}}\right\|_{1}.$
Moreover, GoodDrag faced increased optimization difficulty from this design, due to the larger feature distance compared to the original motion supervision loss. To mitigate this, a smaller step size and more motion supervision steps are used for optimization. This strategy is also applied in DragText.
Appendix C Implementation Details
@@
In Table 2, we list the hyperparameters used for each point-based image editing method [29, 12, 13, 40]. These values are consistently used in both the Baseline and w/ DragText experiments. For a fair comparison, we applied the same hyperparameter values from each respective paper to our experiments. Additionally, we maintained the same text optimization loss across all methods to demonstrate the robustness of our approach.
@@
Appendix D More Qualitative Results
@@
Appendix E Evaluation Metrics
E.1 LPIPS
@@
LPIPS is an appropriate metric for measuring the similarity between two images, emphasizing that image editing should maintain similarity to the original image. However, due to the nature of the drag editing task, the image will inevitably change. Consequently, when dragging is performed successfully, the LPIPS score might worsen. For instance, if an image does not change at all, it would yield an LPIPS score of 0, the best possible score. As shown in Fig. 13, even though we achieved a more desirable image editing outcome, the LPIPS score was lower. Therefore, we propose that LPIPS should not be overly emphasized if the score falls below a certain threshold. To address this issue, we suggest using the product of LPIPS and MD, which are complementary metrics, as a more robust evaluation metric.
@@
E.2 Mean Distance
@@
Mean distance (MD) is computed via DIFT [33]. First, DIFT identifies corresponding points in the edited image that correspond to the handle points in the original image. These identified points are regarded as the final handle points after editing is complete. Then, the mean Euclidean distance between the corresponding point and the target point is calculated. MD is the average value of all point pairs.
We propose that evaluating drag editing using Mean Distance (MD) on certain images in the DragBench dataset is challenging. Some images in DragBench require specific objects to disappear through drag editing as the points move. However, if a specific object disappears, there are no corresponding objects in the edited image, resulting in a significantly high MD value. For instance, in Fig. 14, the handle point and target point indicate that the toothpick should be perfectly inserted into the hamburger. Despite successfully achieving this, DIFT fails to recognize the toothpick, resulting in a higher MD value being calculated. Conversely, there are cases where the MD value is low because the points remain in the same semantic position, but the actual image editing was unsuccessful due to distorted shapes and loss of semantics. While MD is an excellent metric for tasks involving moving feature points of objects, it has certain limitations and challenges when applied to all images in point-based editing tasks.
Appendix F Visual Ablation on the Hyperparameters of Regularization
In Fig. 15, we provide extra visual ablation results to demonstrate how the hyperparameter $\lambda_{\text{text}}$ impacts the regularization process in text optimization. We modified images by adjusting $\lambda_{\text{text}}$ within a range from 0 to 10, which allowed us to control the level of regularization applied during the text optimization phase. When $\lambda_{\text{text}}$ is close to 0, it results in some of the important semantic information being lost. On the other hand, applying an excessively high $\lambda_{\text{text}}$ prevents the optimization of the text embedding from effectively altering the image.

### v2

@@ 8 Conclusion
Figure 10: Limitations of diffusion-based editing methods. (a) The diffusion-based methods sometimes lose important features during the drag optimization process, resulting in content loss, especially when the context prompt lacks specificity. (b) However, if the user provides precise content information, specifying “grapes” as the target to be dragged, DragText successfully performed the drag while preserving the content, unlike the baseline model.
In this study, we introduced DragText, a method emphasizing the critical role of text prompts in point-based image editing with diffusion models. Our findings revealed that static text embeddings hinder the editing process by causing drag halting and a loss of semantic integrity. By optimizing text embeddings in parallel with image embeddings, DragText ensures that text and image embeddings remain strongly coupled, ensuring better drag accuracy and content preservation. Moreover, it integrates seamlessly with existing diffusion-based drag methods, leveraging their strengths while consistently enhancing performance and fidelity in image editing. Our work underlines the importance of considering text-image interaction in point-based editing, providing a novel insight for future developments.
Acknowledgments. This work was supported by IITP RS-2024-00457882 (National AI Research Lab Project), NRF RS-2024-00345806, and NRF RS-2023-00219019 funded by Korean Government (MSIT).
Appendix A Additional Material: Code & Project Page
The code has been submitted as a zip file along with the Supplement. Our results are presented in an easily accessible format on our project page. The link to the project page is as follows: https://micv-yonsei.github.io/dragtext2025/
Appendix B More Details on Prompt Engineering
@@ B.1 Implementation Details
$\hat{\mathbf{c}}^{k}=w^{k}\cdot\hat{\mathbf{c}}^{k}_{\text{org}}+(1-w^{k})\cdot\hat{\mathbf{c}}^{k}_{\text{int}}$
where $\hat{\mathbf{c}}^{k}_{\text{org}}$ is the original text embedding and $\hat{\mathbf{c}}^{k}_{\text{int}}$ is the intention text embedding. At the beginning of the point-based editing, $w^{k}=1$ so $\hat{\mathbf{c}}^{k}=\hat{\mathbf{c}}^{k}_{\text{org}}$. Conversely, as the dragging progresses and handle points approach target points, the weight value $w^{k}$ increases, resulting in a higher proportion of $\hat{\mathbf{c}}^{k}_{\text{int}}$. In this way, as the image is progressively edited, the proportion of the intention text embedding is gradually increased.
B.2 More Qualitative Results for Prompt Engineering
In Fig. 11, we present additional results for prompt engineering. Corresponding results for DragText are also presented to validate the effectiveness of our approach in comparison to prompt engineering. Prompt engineering was found to have little impact on alleviating drag halting. In contrast, DragText dragged handle points closer to target points, compared to the original text prompt, the intention text prompt, and their interpolation.
Figure 11: More qualitative results for prompt engineering (i.e. the intention text, and interpolated intention text). Red points and blue points represent handle points and their target points respectively. Overall, DragText moved the content at the handle point of the User Drag image closer to the target point. Additionally, in (b), it was observed that interpolating the intention text with the original text makes it challenging to maintain the object’s style. In (d), DragText preserved semantics better than other methods while dragging appropriately.
Appendix C More Details on DragText
@@
C.1 Pseudo Code of DragText
@@
Output: Output image $\hat{\mathbf{x}}_{0}$
@@
4:     $\mathcal{F}(\hat{\mathbf{z}}_{t}^{k},\hat{\mathbf{c}}^{k})\leftarrow U_{\theta}(\hat{\mathbf{z}}_{t}^{k};\hat{\mathbf{c}}^{k})$
5:     Update $\hat{\mathbf{z}}_{t}^{k}$ using motion supervision
6:     Update $\hat{\mathbf{c}}^{k}$ using text optimization
7:     $\hat{\mathbf{z}}_{t}^{k+1},\hat{\mathbf{c}}^{k+1}\leftarrow\hat{\mathbf{z}}_{t}^{k},\hat{\mathbf{c}}^{k}$
@@
9:$\hat{\mathbf{z}}_{t},\hat{\mathbf{c}}\leftarrow\hat{\mathbf{z}}_{t}^{K},\hat{\mathbf{c}}^{K}$
@@
11:     $\hat{\mathbf{z}}_{t-1}\leftarrow$ apply denoising to $\hat{\mathbf{z}}_{t}$ conditioned on $\hat{\mathbf{c}}$
12:$\hat{\mathbf{x}}_{0}\leftarrow\hat{\mathbf{z}}_{0}$
@@
C.2 Modifications for Integrate with Other Methods
@@ C.2.1 FreeDrag
$\mathcal{L}_{\text{ms}}=$ $\sum_{i=1}^{n}\sum_{q_{1}}\left\|\mathcal{F}_{q_{1}+d_{i}}(\hat{\mathbf{z}}_{t}^{k},\hat{\mathbf{c}}^{k})-\mathcal{T}_{i}^{k}\right\|_{1}$
$+\lambda_{\text{image}}\left\|\left(\hat{\mathbf{z}}_{t-1}^{k}-\text{sg}(\hat{\mathbf{z}}_{t-1}^{0})\right)\odot(\mathbbm{1}-M_{\text{image}})\right\|_{1}.$
@@
In FreeDrag, if the template feature is poorly learned (category (c)), the point not only reverts to its previous position but also reuses the template feature map without updating its values. Inspired by this approach, our DragText computes the text loss only during the (a) and (b) processes to align with the image. In cases categorized as (c), the text embedding is excluded from the optimization process. Therefore, we define $\alpha_{i}$ as 1 for cases (1) and (2), and 0 for case (3). The text loss is then defined as follows:
$\mathcal{L}_{\text{text}}=$ $\sum_{i=1}^{n}\sum_{q_{1}}\alpha_{i}\left\|\mathcal{F}_{q_{1}+d_{i}}(\hat{\mathbf{z}}_{t}^{k},\hat{\mathbf{c}}^{k})-\mathcal{T}_{i}^{k}\right\|_{1}$
$+\lambda_{\text{text}}\left\|\left(\hat{\mathbf{c}}^{k}-\text{sg}(\mathbf{c}^{0}))\right)\odot M_{\text{text}}\right\|_{1}.$
@@
In DragText, during image optimization, $\hat{\mathbf{c}}^{k}$ does not undergo gradient descent, and during text optimization, the latent vector $\hat{\mathbf{z}}_{t}^{k}$does not undergo gradient descent.
C.2.2 DragNoise
@@
The bottleneck features $\mathbf{s}_{t}$ effectively capture richer noise semantics and efficiently capture most semantics at an early timestep $t$. Thus, DragNoise optimizes the bottleneck feature $\mathbf{s}_{t}$ of the U-Net instead of the latent vector $\mathbf{z}_{t}$ thereby shortening the back-propagation chain. Accordingly, DragText optimizes $\mathbf{s}_{t}$ instead of $\mathbf{z}_{t}$ during the image optimization processes. For each iteration $k$, $\mathbf{\hat{s}}_{t}^{k}$ undergoes a gradient descent step to minimize $\mathcal{L}_{\text{ms}}$:
$\mathbf{\hat{s}}_{t}^{k+1}=\mathbf{\hat{s}}_{t}^{k}-\eta_{\text{ms}}\frac{\partial\mathcal{L}_{\text{ms}}(\mathbf{\hat{s}}_{t}^{k},\hat{\mathbf{c}}^{k})}{\partial\mathbf{\hat{s}}_{t}^{k}}.$
@@
In DragText, neither the latent vector $\mathbf{z}_{t}^{k}$ nor the bottleneck feature $\mathbf{s}_{t}^{k}$ undergoes gradient descent during the text optimization. So the text optimization procedure is not modified in DragNoise.
C.2.3 GoodDrag
@@
GoodDrag alternates between dragging and denoising, introducing periodic corrections to mitigate accumulated errors. This approach is different from traditional diffusion-based dragging methods. They generally execute all drag operations at once before denoising the optimized noisy latent vector $\hat{\mathbf{z}}_{t}$. During the denoising process, which involves sampling images $\hat{\mathbf{x}}_{0}$ from a noisy latent vector $\hat{\mathbf{z}}_{t}$, perturbations from dragging are corrected. However, if the denoising process is performed only after all drag operations are completed, the errors accumulate too significantly to be corrected with high fidelity. To address this, GoodDrag applies one denoising operation after $B$ image optimization and point tracking steps.
@@
For example, the latent vector $\hat{\mathbf{z}}^{k}_{t}$ has been denoised $\lfloor\frac{k}{B}\rfloor$ times, the drag optimization is performed at the timestep $t=T-\frac{k}{B}$. To ensure this process is consistent, the total number of drag steps $K$ should be divisible by $B$. Since DragText performs one text optimization step after one image optimization step, we sequentially repeat the image optimization, text optimization, and point tracking steps $B$ times, and then apply one denoising operation.
@@
$\mathcal{L}_{\text{ms}}=$ $\sum_{i=1}^{n}\sum_{q_{1}}\left\|\mathcal{F}_{q_{1}+4\cdot d_{i}}(\hat{\mathbf{z}}_{t}^{k},\hat{\mathbf{c}}^{k})-\text{sg}(\mathcal{F}_{q_{1}^{0}}(\hat{\mathbf{z}}_{t}^{0},\hat{\mathbf{c}}^{0}))\right\|_{1}$
$+\lambda_{\text{image}}\left\|\left(\hat{\mathbf{z}}_{t-1}^{k}-\text{sg}(\hat{\mathbf{z}}_{t-1}^{0})\right)\odot(\mathbbm{1}-M_{\text{image}})\right\|_{1},$
where $q_{1}=\Omega(h_{i}^{k},r_{1})$, and $q_{1}^{0}$ describes the square region centered at the original handle point $h_{i}^{0}$. And, drag operations per denoising step $B=10$. Similarly, DragText ensures the handle point $h^{k}_{i}$ remains consistent with the original point $h^{0}_{i}$ during text optimization:
$\mathcal{L}_{\text{text}}=$ $\sum_{i=1}^{n}\sum_{q_{1}}\left\|\mathcal{F}_{q_{1}+4\cdot d_{i}}(\hat{\mathbf{z}}_{t}^{k},\hat{\mathbf{c}}^{k})-\text{sg}(\mathcal{F}_{q_{1}^{0}}(\hat{\mathbf{z}}_{t}^{0},\hat{\mathbf{c}}^{0}))\right\|_{1}$
$+\lambda_{\text{text}}\left\|\left(\hat{\mathbf{c}}^{k}-\text{sg}(\mathbf{c}^{0}))\right)\odot M_{\text{text}}\right\|_{1}.$
Additionally, GoodDrag faced increased optimization difficulty from this design, due to the larger feature distance compared to the original motion supervision loss. To mitigate this, a smaller step size and more motion supervision steps are used for optimization. This strategy is also applied in DragText.
Appendix D Implementation Details
@@
In Table 2, we listed the hyperparameters used for each point-based image editing method [31, 13, 14, 42]. These values were consistently used in both the Baseline and w/ DragText experiments. For a fair comparison, we applied the same hyperparameter values from the respective paper to our experiments. Additionally, we maintained the same text optimization loss across all methods to demonstrate the robustness of our approach.
@@
Appendix E More Qualitative Results
@@
Appendix F Evaluation Metrics
F.1 LPIPS
@@
LPIPS is an appropriate metric for measuring the similarity between two images, emphasizing that image editing should maintain similarity to the original image. However, due to the nature of the drag editing task, the image will inevitably change. Consequently, even when dragging is performed successfully, the LPIPS score might worsen. For instance, if an image does not change at all, it would yield an LPIPS score of 0, the best possible score. As shown in Fig. 13, even though we achieved a more desirable image editing outcome, the LPIPS score was lower. Therefore, we propose that LPIPS should not be overly emphasized if the score falls below a certain threshold. To address this issue, we suggest using the product of LPIPS and MD, which are complementary metrics, as a more robust evaluation metric.
F.2 Mean Distance
@@
Mean distance (MD) is computed via DIFT [35]. First, DIFT identifies corresponding points in the edited image that correspond to the handle points in the original image. These identified points are regarded as the final handle points after editing is complete. Then, the mean Euclidean distance between the corresponding point and the target point is calculated. MD is the average value of all handle-target point pairs.
We propose that evaluating drag editing using Mean Distance (MD) on certain images in the DragBench dataset is challenging. Some images in DragBench require specific objects to disappear through drag editing as the points move. However, if a specific object disappears, there would be no corresponding objects in the edited image, resulting in a significantly high MD value. For instance, in Fig. 14, the handle point and target point indicate that the toothpick should be perfectly inserted into the hamburger. Despite successfully achieving this, DIFT fails to recognize the toothpick, resulting in a higher MD value being calculated. Conversely, there are cases where the MD value is low because the points remain in the same semantic position, but the actual image editing was unsuccessful due to distorted shapes and loss of semantics. While MD is an excellent metric for tasks involving moving feature points of objects, it has certain limitations and challenges when applied to all images in point-based editing tasks.
Appendix G Visual Ablation on the Hyperparameters of Regularization
In Fig. 15, we provide extra visual ablation results to demonstrate how the hyperparameter $\lambda_{\text{text}}$ impacts the regularization process in text optimization. We modified images by adjusting $\lambda_{\text{text}}$ within a range from 0 to 10, which allowed us to control the level of regularization applied during the text optimization phase. When $\lambda_{\text{text}}$ is close to 0, it results in some of the important semantic information being lost. On the other hand, applying an excessively high $\lambda_{\text{text}}$, prevents the optimization of the text embedding from effectively altering the image.